OpenStudy (bbb911):

Find the total-profit function for the given total-cost and total-revenue functions. C(x) = 58x + 912,000 R(x) = 260x

ganeshie8 (ganeshie8):

`Cost` is the cost of manufacturing something `Revenue` comes from sales so, Profit = Revenue - Cost

ganeshie8 (ganeshie8):

just subtract C(x) from R(x)

ganeshie8 (ganeshie8):

Profit = R(x) - C(x) = 260x - (58x + 912,000) = ?